# InceptionMamba: An Efficient Hybrid Network with Large Band Convolution and Bottleneck Mamba

### v2

@@ 4.1 Image classification on ImageNet
For classification, the public ImageNet-1k [18, 19] dataset is involved in our evaluations. Our model is trained from scratch at 224$\times$224 resolution for 300 epochs using the AdamW optimizer with a weight decay of 5e-2, initial learning rate of 1e-3, and batch size of 512. The training strategy incorporates a cosine scheduler with 10 warmup epochs, Label Smoothing of 0.1, stochastic depth, and RandAugment. For fairness, we adopt conventional data augmentation methods including Mixup, random erasing, and AutoAugment following [20]. Numerous advanced state-of-the-art backbone networks are involved in our comparative studies, including Vision-RWKV [21], Vision Mamba [15], Vision LSTM [22], PlainMamba [23], EfficientVMamba [24], VMamba [9], InceptionNeXt [4], MambaVision [8], QuardMamba [25], PeLK [26], MambaOut [16] and RDNet [27]. Parallel training is performed with eight NVIDIA RTX 4090 GPUs in all our experiments.
@@ 4.2 Object detection &\&& instance segmentation on COCO
For object detection and instance segmentation, we have conducted experiments on MS-COCO which contains 118k training images and 5k validation images [28]. The detection framework is utilized as Mask R-CNN [29] with our InceptionMamba backbone, and training the detector is initialized with the model pretrained on the ImageNet. In implementation, we follow the standard 1$\times$ training protocol (12 epochs), and adopt multi-scale data augmentation. To maintain the original aspect ratio, the input image is resized such that the shorter side is fixed at 800 pixels while the longer side does not exceed 1333 pixels. Using the AdamW optimizer, initial learning rate is set to 0.0002, whilst weight attenuation coefficient is 0.05. In comparative studies, competing architectures range from CNN and Transformer like ConvNeXt [3] and Swin-Transformer [30] to Mamba-based networks, e.g., EfficientVMamba and MambaOut.
@@ 4.3 Semantic segmentation on ADE20K
In addition to the above-mentioned evaluations, we perform semantic segmentation experiments on public ADE20K dataset [34] using UperNet architecture [35]. This dataset has 150 semantic categories with 20,000 training images and 2,000 validation images. Initialized with the pretrained model similar to the detection experiments, we follow [16] to adopt a protocol of 160,000 iterations in our training process, and utilize the AdamW optimizer with an initial learning rate set to 0.00012 and a batch size of 16.
@@ 4.4 Ablation studies
To investigate the impact of different branch allocation ratios on model performance, in the structure without the Bottleneck Mamba, we conducted ablation experiments comparing three distinct channel allocation strategies. The experimental results demonstrate that when employing a ratio of [0.125, 0.125, 0.75], the model achieves 81.9% Top-1 accuracy with 21.4M parameters and 3.35G FLOPs computational cost, exhibiting the optimal balance between accuracy and efficiency. In contrast, increasing the ratio of the second branch to 0.25 ([0.125, 0.25, 0.625]) leads to a slight parameter increase to 21.5M and a marginal performance drop (81.7%), indicating that the asymmetric convolutional branches are sensitive to channel variations. A balanced allocation scheme ([0.25, 0.25, 0.5]) maintains 81.9% accuracy with comparable computational cost (3.38G FLOPs), verifying the robustness of the module design to different allocation ratios. After comprehensive evaluation, we select [0.125, 0.125, 0.75] as the final configuration, as it achieves the best trade-off between computational efficiency and model performance.
Table 5: Comparison Results of different Branch Ratio
@@
Local mixer
We conducted comprehensive ablation studies by replacing our token mixer with two established designs: (1) ConvNeXt’s [3] depthwise convolutions and (2) InceptionNeXt’s [4] decomposed convolution. We also replaced band convolution with strip convolution for testing. Experimental results demonstrate that our proposed token mixer achieves superior performance while maintaining identical model complexity.
Table 6: Comparison Results of Different Local Mixers
Local mixer Params(M) FLOPs(G) Top-1(%)
@@
The ablation study demonstrates that integrating the Bottleneck structure with SS2D [9] maintains identical model accuracy while significantly improving efficiency. Compared to the baseline SS2D [9], the Bottleneck-enhanced version achieves the same Top-1 performance while reducing both parameters and computational costs by over 10%. Additionally, the throughput increases substantially, highlighting the Bottleneck’s ability to compress feature dimensions effectively without sacrificing performance.
Table 7: Comparsion Results of SS2D and Bottleneck + SS2D
@@
Module in Bottleneck
The ablation experiments systematically evaluate different bottleneck-integrated architectures, with Bottleneck + SS2D emerging as the optimal configuration. Compared to the baseline without Mamba Bottleneck (21.4M Params, 3.3G FLOPs, 81.9% Top-1), all bottleneck variants improve accuracy, but Bottleneck + SS2D achieves the highest Top-1 accuracy (83.1%) while maintaining competitive efficiency—reducing FLOPs by 13% compared to Bottleneck + attention (4.0G vs. 4.6G) and doubling the throughput (1362 imgs/s vs. 1305 imgs/s). Notably, while Bottleneck + attention (using EfficientViT’s [36] LiteMLA linear attention) slightly improves accuracy (82.3%), its computational cost is significantly higher, highlighting SS2D’s superior balance between performance and efficiency. Simpler designs like Bottleneck + GELU or Bottleneck + DWConv7×7 achieve marginal accuracy gains (82.1%) but fail to match SS2D’s global feature capture capability. These results underscore that SS2D, when combined with bottleneck structures, optimally leverages hardware-friendly operations (e.g., linear complexity scans) to enhance both accuracy and speed, aligning with the efficiency goals of modern lightweight architectures.
Table 8: Comparsion Results of different module in Bottleneck
@@
More InceptionMamba Heatmap
The ablation study of the heatmap shows that InceptionMamba outperforms ConvNeXt [3] and Inceptionnext [4]. For the input images of a dog, InceptionMamba produces heatmaps that more comprehensively and accurately highlight object regions. It effectively highlights the dog’s limbs. This indicates that the InceptionMamba architecture, likely due to its unique Mamba-based temporal processing combined with Inception-like multi-scale feature fusion, improves global context modeling. The enhanced heatmap coherence implies that InceptionMamba captures richer spatial relationships, presenting it as a potential state-of-the-art lightweight architecture with qualitative superiority in visual feature attribution, which also complements previous ablation results regarding efficiency-accuracy trade-off.
Input
ConvNeXt
Inceptionnext
InceptionMamba
Figure 5: Comparison of heatmaps generated by different models.
@@ 5 Conclusion
In this paper, we introduce a novel hierarchical vision architecture named InceptionMamba through the lens of multi-scale spatial global interaction. Built upon the InceptionNeXt [4] framework, we present a modular architecture that leverages Large Band Depthwise Convolution as the Local Mixer for extracting local features and adopts Mamba [13, 14] Bottleneck as Global Mixer for global modeling, thereby enabling the network to capture local details and long-range semantic dependencies across hierarchical token spaces. Extensive experiments demonstrate the consistent superiority of InceptionMamba in terms of both accuracy and computational efficiency compared to representative ConvNets (e.g., ConvNeXt [3]) and Transformer (e.g., Swin [30]) on benchmarks spanning image classification and several downstream vision tasks. We hope our work can inspire the exploration of efficient multi-scale-global interaction paradigms in deep architecture design and facilitate the development of lightweight yet powerful vision backbones.

### v3

@@ 4.1 Image classification on ImageNet
For classification, the public ImageNet-1k [28, 29] dataset is involved in our evaluations. Our model is trained from scratch at 224$\times$224 resolution for 300 epochs using the AdamW [30] optimizer with a weight decay of 5e-2, initial learning rate of 1e-3, and batch size of 512. The training strategy incorporates a cosine scheduler with 10 warmup epochs, Label Smoothing of 0.1, stochastic depth, and RandAugment. For fairness, we adopt conventional data augmentation methods including Mixup, random erasing, and AutoAugment following [31]. Numerous advanced state-of-the-art backbone networks are involved in our comparative studies, including Vision-RWKV [32], Vision Mamba, Vision LSTM [33], PlainMamba [34], EfficientVMamba [35], VMamba, InceptionNeXt, MambaVision, QuardMamba [36], PeLK [37], MambaOut and RDNet [38]. Parallel training is performed with eight NVIDIA RTX 4090 GPUs in all our experiments.
@@ 4.2 Object detection &\&& instance segmentation on COCO
For object detection and instance segmentation, we have conducted experiments on MS-COCO [39] which contains 118k training images and 5k validation images. The detection framework is utilized as Mask R-CNN [40] with our InceptionMamba backbone, and training the detector is initialized with the model pretrained on the ImageNet. In implementation, we follow the standard 1$\times$ training protocol (12 epochs), and adopt multi-scale data augmentation. To maintain the original aspect ratio, the input image is resized such that the shorter side is fixed at 800 pixels while the longer side does not exceed 1333 pixels. Using the AdamW optimizer, initial learning rate is set to 0.0002, whilst weight attenuation coefficient is 0.05. In comparative studies, competing architectures range from CNN and Transformer like ConvNeXt and Swin-Transformer to Mamba-based networks, e.g., EfficientVMamba  and MambaOut.
@@ 4.3 Semantic segmentation on ADE20K
In addition to the above-mentioned evaluations, we perform semantic segmentation experiments on public ADE20K dataset [45] using UperNet architecture [46, 47]. This dataset has 150 semantic categories with 20,000 training images and 2,000 validation images. Initialized with the pretrained model similar to the detection experiments, we follow [26] to adopt a protocol of 160,000 iterations in our training process, and utilize the AdamW optimizer with an initial learning rate set to 0.00012 and a batch size of 16.
@@ 4.4 Ablation studies
To investigate the impact of different branch allocation ratios on model performance within ConvMixer, we have conducted ablation experiments comparing three different channel allocation strategies in our model excluding GlobalMixer module. The experimental results as shown in Table 5 demonstrate that the best 81.9% Top-1 accuracy is reported when employing a ratio of [0.125, 0.125, 0.75] with 21.4M parameters and 3.35G FLOPs, exhibiting the optimal balance between accuracy and efficiency. In contrast, increasing the ratio of the second branch to 0.25 ([0.125, 0.25, 0.625]) leads to a slight parameter increase to 21.5M and a marginal performance drop (81.7%), indicating that the asymmetric convolutional branches are sensitive to channel variations. A balanced allocation scheme ([0.25, 0.25, 0.5]) maintains 81.9% accuracy with comparable computational cost (3.38G FLOPs), verifying the robustness of the module design to different allocation ratios. After comprehensive evaluation, the branch ratio is empirically set as [0.125, 0.125, 0.75] in our remaining experiments, as it achieves the best accuracy-efficiency trade-off.
Table 5: Comparison Results of different settings of Branch Ratios (square conv./Band conv./Identity Mapping) within ConvMixer. The GlobalMixer is excluded from our InceptionMamba.
@@
ConvMixer
To delve into the ConvMixer within our InceptionMamba, we have conducted comprehensive ablation studies by comparing our ConvMixer with two different designs: Depthwise convolutions within ConvNeXt and decomposed convolutions within InceptionNeXt. We also replaced band convolution with strip convolution for comparison. As demonstrated in Table 6, our proposed ConvMixer achieves superior performance while maintaining identical model complexity.
Table 6: Comparison Results of Different structures within our ConvMixer.
ConvMixer Params(M) FLOPs(G) Top-1(%)
@@
Moreover, we explore the benefits of bottleneck structure within our GlobalMixer. As shown in Table 7, integrating the bottleneck with SS2D maintains model accuracy while significantly improving efficiency. Compared to the baseline SS2D, introducing the bottleneck achieves the same Top-1 accuracy while reducing both parameters and computational costs by over 10%. Additionally, the throughput increases substantially, highlighting the ability of the bottleneck in efficiently compressing feature dimensions without sacrificing model performance as aforementioned.
Table 7: Comparison of SS2D (Mamba) and Bottleneck + SS2D (GlobalMixer).
@@
GlobalMixer
To further investigate the GlobalMixer within our framework, we have systematically evaluated different architectures involving the bottleneck as illustrated in Table 8. Compared to the baseline without Mamba Bottleneck, all the other variants involving bottleneck improve Top-1 accuracy. In particular, our GlobalMixer achieves the highest 83.1% Top-1 accuracy, while maintaining competitive efficiency by reducing FLOPs by 13% compared to the structure of Bottleneck $+$ attention (LiteMLA linear attention in EfficientViT [48]) and promoting the throughput from 1305 imgs/s to 1362 imgs/s. Notably, while Bottleneck $+$ attention slightly improves accuracy, it leads to significantly increased computational cost. This contrast suggests the advantage of SS2D in balancing between performance and efficiency. Simpler designs like Bottleneck $+$ GELU or Bottleneck $+$ DWConv7×7 achieve marginal performance gains but fail to rival SS2D in terms of global modeling capability. These results demonstrate that our GlobalMixer incorporating SS2D and bottleneck optimally leverages hardware-friendly operations (e.g., linear complexity scans) to achieve accuracy-speed trade-off, thereby well aligning with the efficiency goal of modern lightweight architectures.
Table 8: Comparison of different modules involving Bottleneck structure.
@@
More Visualization Results
To qualitatively demonstrate the advantages of the proposed InceptionMamba, we intuitively compare our method with the classic ConvNeXt and InceptionNeXt frameworks. As illustrated in Fig. 6, given an input “dog” image, InceptionMamba produces CAMs that can capture object-aware regions more comprehensively and accurately, leading to high responses to the regions of the dog’s limbs and head. In contrast, for the input “cat” image, ConvNeXt mainly focuses on the face of the cat, while InceptionNeXt pays attention to the cat’s legs area. In contrast, our InceptionMamba can capture key regions including the cat’s head and body area that are highly semantically relevant to the “cat” class. This indicates that our InceptionMamba architecture can improve global context modeling capacity while inherits multi-scale feature fusion from Inception-like structures. Besides, the qualitative comparisons demonstrate that InceptionMamba can capture richer spatial relationships, manifesting itself as a potential state-of-the-art lightweight architecture. This also complements previous ablation results regarding the advantage of our method in efficiency-accuracy trade-off.
Figure 6: Comparison of CAMs generated from different mainstream architectures. Our InceptionMamba is capable of characterizing key semantic-aware regions, which is beneficial for a variety of downstream vision tasks.
@@ 5 Conclusion
In this study, we propose a hybrid backbone architecture termed InceptionMamba. Inheriting efficient parallel structure from InceptionNeXt framework, we leverage large band depthwise convolutions within our ConvMixer for capturing local cues and develop bottleneck Mamba for channel interaction, such that long-range dependencies modeling capability can be significantly improved. Extensive experiments demonstrate the superiority of InceptionMamba in terms of both accuracy and computational efficiency compared to existing mainstream CNNs and ViT architectures on various public benchmarks spanning image classification and several downstream vision tasks. We hope our work can inspire the exploration of efficient multi-scale global interaction paradigms in deep architecture design and facilitate the development of lightweight yet powerful vision backbones.
Acknowledgment
This work was supported by the National Natural Science Foundation of China under Grant 62173186, 62076134, 62276061, and the Key R & D Program of Zhejiang Province (2024C04056(CSJ)).